OpenStudy (anonymous):

Limit X-->0 cot(4x)/csc(3x) ?

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

\[\frac {\cot(4x)}{cosec(3x)}=\cos(4x)\frac{\sin(3x)}{\sin(4x)}\]\[\lim_{x \rightarrow 0} \cos(4x)\frac {\sin(3x)}{3x}\frac{4x}{\sin(4x)}\frac{3}{4}=\frac{3}{4}\]

OpenStudy (anonymous):

hope you got it.

OpenStudy (anonymous):

it's very clear, THANKS